A 15000-kg helicopter is lifting a 4500-kg car with an upward acceleration of 1.4 m/s^2.calculate (a)the vertical force of air exerts on the helicopter blades and (b) the tension in the upper supporting caple

Grade:12th pass

1 Answers

Prateek Mishra
3 years ago
  • Suppose we have just to maintain the halicopter-car system in air 
  • then the wings had to apply the equal force as the gravity is applying 
  • but we have to go up by an acceleration of 1.4m/s2  so the effective gravity wil be (9.8+1.4)
  • Fnet=(15000+4500)(9.8+1.4)=you may solve
  •  samewise the tension in the wire(pulling the car up)
  • tnet=mass of the car*geff=4500*10.2=